(1 point) In each part, we have given the significance level and the P-value for a hypothesis test. For each case determine if the null hypothesis should be rejected. Write reject or do not reject (without quotations). a)a 0.06, P 0.06 answer. (b) α 0.07,P-0.06 answer (c) α 0.01, P. 006 answer Note: You can earn partial credit on this problem. Preview My Answers Submit Answers You have attempted this problem 0 times. You have unlimited attempts remaining
Answer #1

The steps involved in using the p-value approach to conduct any hypothesis test are :

1. Specify the null and the alternative hypothesis.

2. Using the sample data and assuming that the null hypothesis is true, calculate the value of test statistics. For eg: For testing of population mean of a population from normal distribution , we use the test statistic t=(\bar{x}-\mu )/(s*\sqrt{n}) ~ t(n-1), always remember this test is used when population variance is unknown ,otherwise we use z test.

3. Using the known distribution of the test statistics, calculate the p-value.

4. Let \alpha be the level of significance.

Decision : If p-value is less than or equal to \alpha , reject the null hypothesis.

If p-value is greater than \alpha , do not reject the null hypothesis.

Answers:

1: \alpha =0.06 and p=0.06

Since, p<= \alpha , we reject the null hypothesis.

2: \alpha =0.07 and p=0.06

Since, p< \alpha , we reject the null hypothesis.

3: \alpha =0.01 and p=0.06

Since, p> \alpha , we do not reject the null hypothesis.